# Relatively big projective modules and their applications to direct sum decompositions

### v1

@@ 6.2. Locally semiperfect algebras over $h$-local domains
Let $\Lambda$ be an algebra over a commutative ring $R$. We say that $\Lambda$ is locally semiperfect if $\Lambda_{\mathfrak{m}}$ is semiperfect for every maximal ideal $\mathfrak{m}$ of $R$.
In this section, it is important to keep in mind that we are using the conventions of Notation 5.1.
The following example explains a particular instance of the appearance of locally semiperfect algebras in which we will be particularly interested in Section 7.
@@
Proof.  Let $I$ be a non-zero two-sided ideal of $\Lambda$ which is the trace of a projective right $\Lambda$-module. Since $\Lambda$ is torsion-free, $I_{\mathfrak{m}}\neq\{0\}$ for every maximal ideal $\mathfrak{m}$ of $R$. By Lemma 5.3 and Lemma 6.2, $I_{\mathfrak{m}}$ is also the trace of a finitely generated projective right $\Lambda_{\mathfrak{m}}$-module.
Let $S$ denote a simple right $\Lambda_{Q}$-module, and assume that $\Lambda_{Q}\cong S^{k}$ as right $\Lambda_{Q}$-modules. Since $\Lambda_{Q}$ is simple artinian and $I$ is non-zero, by Lemma 5.8, there exists a finite set $\mathcal{M}$ of maximal ideals of $R$ such that $I_{\mathfrak{m}}=\Lambda_{\mathfrak{m}}$ if $\mathfrak{m}\not\in\mathcal{M}$. For any maximal ideal $\mathfrak{m}\in\mathcal{M}$, let $P(\mathfrak{m})$ be a finitely generated projective right $\Lambda_{\mathfrak{m}}$-module with trace ideal $I_{\mathfrak{m}}$. Then $P(\mathfrak{m})_{Q}\cong S^{t_{\mathfrak{m}}}$ for a suitable positive integer $t_{\mathfrak{m}}\geq 1$. Let $t=\mathrm{lcm}\{t_{\mathfrak{m}}\mid\mathfrak{m}\in\mathcal{M}\}$, write $t=\ell_{\mathfrak{m}}\cdot t_{\mathfrak{m}}$, and let $P^{\prime}(\mathfrak{m}):=P(\mathfrak{m})^{t\ell_{\mathfrak{m}}}$ for each maximal ideal $\mathfrak{m}$ of $\mathcal{M}$. For each $\mathfrak{m}\notin\mathcal{M}$, take $P^{\prime}(\mathfrak{m})=\Lambda_{\mathfrak{m}}^{t}$. Note that $P^{\prime}(\mathfrak{m})_{Q}\cong S^{tk}\cong\Lambda_{Q}^{t}$, and $\operatorname{Tr}_{\Lambda_{\mathfrak{m}}}(P^{\prime}(\mathfrak{m}))=I_{\mathfrak{m}}$ for every maximal ideal $\mathfrak{m}$ of $R$. By Package Deal Theorem 5.12 applied the package $\{P^{\prime}(\mathfrak{m})\mid\mathfrak{m}\in\operatorname{mSpec}R\}$ there exists a finitely generated projective right $\Lambda$-module $P$ such that $P_{\mathfrak{m}}\cong P^{\prime}(\mathfrak{m})$ for any maximal ideal $\mathfrak{m}$ of $R$. By Lemma 5.3, $\operatorname{Tr}_{\Lambda}(P)_{\mathfrak{m}}=\operatorname{Tr}_{\Lambda_{\mathfrak{m}}}(P^{\prime}(\mathfrak{m}))=I_{\mathfrak{m}}$ for any maximal ideal $\mathfrak{m}$ of $R$. This implies that $I=\operatorname{Tr}_{\Lambda}(P)+I=\operatorname{Tr}_{\Lambda}(P)$.
A similar argument gives the conclusion for the traces of projective left $\Lambda$-modules because of the symmetry of the hypotheses.
The equality $\mathcal{T}_{r}(\Lambda)=\mathcal{T}_{\ell}(\Lambda)$ follows because of the canonical bijection between finitely generated projective right and left $\Lambda$-modules induced by the $\Lambda$-dual, and because the trace of a finitely generated projective module coincides with the trace of its $\Lambda$-dual, cf. Remark 2.9.
@@
Let $R$ be a commutative domain of finite character with field of fractions $Q$. Let $\Lambda$ be a locally semiperfect torsion-free $R$-algebra such that $\Lambda_{Q}$ is simple artinian; so that there exist $k$ and a division ring D such that $\Lambda_{Q}\cong M_{k}(D)$. Then for almost all maximal ideals $\mathfrak{m}$ of $R$ the ring $\Lambda_{\mathfrak{m}}$ has the following properties:
@@
    $\Lambda_{\mathfrak{m}}$ has exactly one indecomposable projective up to an isomorphism. Moreover, $\Lambda_{\mathfrak{m}}$ is isomorphic to a ring of $k\times k$ matrices over a local ring $T_{\mathfrak{m}}$ and $T_{\mathfrak{m}}\otimes_{R}Q\cong D$.
@@
    If $U$ and $V$ are projective $\Lambda_{\mathfrak{m}}$-modules then $U\cong V$ if and only if $U_{Q}\cong V_{Q}$.
Proof.  Let $S$ be a simple $\Lambda_{Q}$-module and assume that $\Lambda_{Q}\cong S^{k}$. Let $X$ be a finitely generated $\Lambda$-module, which is torsion-free as an $R$-module, and such that $X_{Q}\cong S$. By Lemma 5.7, $X_{\mathfrak{m}}^{k}$ is free for almost all maximal ideals. If $\mathfrak{m}$ is such a maximal ideal, then $X_{\mathfrak{m}}^{k}\cong\Lambda_{\mathfrak{m}}$ gives an indecomposable decomposition of the semiperfect ring $\Lambda_{\mathfrak{m}}$. Hence $X_{\mathfrak{m}}$ is the only indecomposable projective $\Lambda_{\mathfrak{m}}$-module up to an isomorphism. As $\Lambda_{\mathfrak{m}}$ is semiperfect, the endomorphism ring of $X_{\mathfrak{m}}$ is a local ring that we will denote by $T_{\mathfrak{m}}$. Hence $\Lambda_{\mathfrak{m}}\cong\mathrm{End}_{\Lambda_{\mathfrak{m}}}(X_{\mathfrak{m}}^{k})\cong M_{k}(T_{\mathfrak{m}}).$ By Lemma 5.2, $T_{\mathfrak{m}}\otimes_{R}Q\cong D$.
@@
Let $R$ be an $h$-local domain with field of fractions $Q$, and let $\Lambda$ be a locally semiperfect torsion-free $R$-algebra such that $\Lambda_{Q}$ is simple artinian. In view of Lemma 6.8, the maximum spectrum of $R$ can be written as a disjoint union of two sets $A$ and $B$ such that for any $\mathfrak{m}\in A$, $\Lambda_{\mathfrak{m}}$ is a ring of matrices over a local ring and $B$ is a finite, non-empty set. In particular, for $\mathfrak{m}\in A$, $\mathcal{T}_{r}(\Lambda_{\mathfrak{m}})=\mathcal{T}_{\ell}(\Lambda_{\mathfrak{m}})=\{0,\Lambda_{\mathfrak{m}}\}$.
In view of the Package Deal Theorem for Traces 5.10, there is a bijective correspondence
@@
Let $R$ be an $h$-local domain with field of fractions $Q$, and let $\Lambda$ be a locally semiperfect torsion-free $R$-algebra such that $\Lambda_{Q}$ is simple artinian. Then $\mathcal{T}(\Lambda)$ is finite. Moreover, if $P$ is a finitely generated projective $\Lambda$-module, then $I=\operatorname{Tr}_{\Lambda}(P)$ is a minimal nonzero element of $\mathcal{T}(\Lambda)$ if and only if $P_{\mathfrak{m}}$ is a finite power of an indecomposable projective $\Lambda_{\mathfrak{m}}$-module for any $\mathfrak{m}\in\operatorname{mSpec}R$.
Proof.  Finiteness of ${\mathcal{T}}(\Lambda)$ follows from Remark 6.9. Also note that the correspondence of the remark preserves inclusions, so $\operatorname{Tr}_{\Lambda}(P)$ is a minimal element of ${\mathcal{T}}(\Lambda)\setminus\{0\}$ if and only if $\operatorname{Tr}_{\Lambda_{\mathfrak{m}}}(P_{\mathfrak{m}})$ is a minimal element of ${\mathcal{T}}(\Lambda_{\mathfrak{m}})\setminus\{0\}$ for any $\mathfrak{m}\in\operatorname{mSpec}(R)$ if and only if $P_{\mathfrak{m}}$ is a finite power of an indecomposable projective $\Lambda_{\mathfrak{m}}$-module.
@@
  1. (1)
    $P_{\Lambda}$ is finitely generated if and only if $P_{Q}$ is a finitely generated $\Lambda_{Q}$-module if and only if there exists a maximal ideal $\mathfrak{m}$ of $R$ such that $P_{\mathfrak{m}}$ is a finitely generated $\Lambda_{\mathfrak{m}}$-module.
  2. (2)
    The set $\mathcal{C}=\{J\unlhd\Lambda\mid P/PJ\mbox{ is finitely generated}\}$ has a minimal element which is the trace of a finitely generated, projective right $\Lambda$-module. If $I$ is such ideal then it is characterized by the property that, for any maximal ideal $\mathfrak{m}$ of $R$, $I_{\mathfrak{m}}$ is the minimal element of $\mathcal{C}_{\mathfrak{m}}=\{J\unlhd\Lambda\mid P_{\mathfrak{m}}/P_{\mathfrak{m}}J\mbox{ is finitely generated}\}$.
@@
$(1)$ We claim that if $P_{Q}\cong S^{n}$ for some $n\in\mathbb{N}$, then $P$ is finitely generated. Indeed, by hypothesis, there exists a finitely generated $\Lambda$-submodule $P_{0}$ of $P$ such that $(P_{0})_{Q}=P_{Q}$. Then we may assume that we have, $P_{0}\leq P\leq\Lambda^{(\omega)}=P\oplus P^{\prime}$, and for any $p\in P$ there exists $0\neq d\in R$ such that $pd\in P_{0}$. Since $\Lambda$ is torsion-free as an $R$-module, this implies that there exists a finite subset $F$ of $\omega$ such that $P_{0}\leq P\subseteq\Lambda^{F}$. Hence $\Lambda^{F}=P\oplus(P^{\prime}\cap\Lambda^{F})$ so $P$ is finitely generated.
This proves statement $(1)$.
$(2)$ In the proof, we keep the notation of Remark 6.9 of the partition of the maximal spectrum of $R$.
If $P$ is finitely generated, the statement is clear. Assume that $P$ is a countably generated, projective, right $\Lambda$-module that is not finitely generated. By $(1)$, $P_{Q}\cong S^{(\omega)}$. Therefore, there are two possible situations with the localizations of $P$ at maximal ideals of $R$:
  • (i)
    $P_{\mathfrak{m}}$ is a free $\Lambda_{\mathfrak{m}}$-module of infinite rank for any $\mathfrak{m}\in A$; so $\Lambda_{\mathfrak{m}}$ is the only element of the set $\mathcal{C}_{\mathfrak{m}}=\{J\mbox{ ideal of $\Lambda_{\mathfrak{m}}$}\mid P_{\mathfrak{m}}/P_{\mathfrak{m}}J\mbox{ is finitely generated}\}$.
  • (ii)
    $P_{\mathfrak{n}}$ is an infinitely generated projective module over the semiperfect ring $\Lambda_{\mathfrak{n}}$ for any $\mathfrak{n}\in B$; by Proposition 6.3, there exists an ideal $I(\mathfrak{n})\neq\{0\}$, minimal element of the set $\mathcal{C}_{\mathfrak{n}}=\{J\mbox{ ideal of $\Lambda_{\mathfrak{n}}$}\mid P_{\mathfrak{n}}/P_{\mathfrak{n}}J\mbox{ is finitely generated}\}$. In addition $I(\mathfrak{n})$ is the trace of a finitely generated projective right $\Lambda_{\mathfrak{n}}$-module.
By the Package Deal Theorem 5.10, there exists $I$ ideal of $\Lambda$ such that $I_{\mathfrak{m}}=\Lambda_{\mathfrak{m}}$ for any $\mathfrak{m}\in A$ and $I_{\mathfrak{n}}=I(\mathfrak{n})$ for any $\mathfrak{n}\in B$. In addition, $I$ is the trace of a (finitely generated, by Lemma 6.6) projective right $\Lambda$-module. We claim that $I$ is the minimal element of $\mathcal{C}$.
@@
which is finitely generated. Therefore, $I\in\mathcal{C}$.
Let $I^{\prime}\in\mathcal{C}$ be such that $I^{\prime}\subseteq I$. Then, because $\Lambda$ is torsion-free as an $R$-module, for any maximal ideal $\mathfrak{m}$ of $R$, $I^{\prime}_{\mathfrak{m}}\in\mathcal{C}_{\mathfrak{m}}$ and $(I^{\prime})_{\mathfrak{m}}\subseteq I_{\mathfrak{m}}=I(\mathfrak{m})$. The minimality in the choice of $I(\mathfrak{m})$ in $(i)$ and $(ii)$, implies that $(I^{\prime})_{\mathfrak{m}}=I_{\mathfrak{m}}=I(\mathfrak{m})$ for any maximal ideal $\mathfrak{m}$ of $R$. Hence, $I^{\prime}=I$. This shows that $I$ is the minimal element of $\mathcal{C}$.
Notice that the proof also shows that, for any maximal ideal $\mathfrak{m}$ of $R$, $I_{\mathfrak{m}}$ is the minimal element of $\mathcal{C}_{\mathfrak{m}}$, which is in $\mathcal{T}(\Lambda_{\mathfrak{m}})$ Proposition 6.3. By the Package Deal Theorem 5.10, $I$ is uniquely determined by its localizations.
@@
Let $R$ be a commutative domain with field of fractions $Q$, and let $\Lambda$ be a torsion-free $R$-algebra. Let $P$ be a countably generated projective $\Lambda$-module, $I\in\mathcal{T}(\Lambda)$, and $V$ a finitely generated $\Lambda$-module which is torsion-free as an $R$-module. Let $\mathfrak{m}$ be a prime ideal of $R$ such that $V_{\mathfrak{m}}=V_{\mathfrak{m}}I_{\mathfrak{m}}$ and $P_{\mathfrak{m}}=\bigoplus_{i\in\mathbb{N}}U_{i}$ with $U_{i}$ finitely generated $\Lambda_{\mathfrak{m}}$-modules such that infinitely many of these $U_{i}$ are isomorphic to a $\Lambda_{\mathfrak{m}}$-module $U$ whose trace ideal contains $I_{\mathfrak{m}}$. Then for every finitely generated submodule $P_{1}$ of $P$, there exists a homomorphism $f\in\operatorname{Hom}_{\Lambda}(P,V)$ such that $f(P_{1})=0$ and $f_{\mathfrak{m}}$ is onto.

### v2

@@ 6.2. Locally semiperfect algebras over $h$-local domains
Let $\Lambda$ be an algebra over a commutative domain $R$. We say that $\Lambda$ is locally semiperfect if $\Lambda_{\mathfrak{m}}$ is semiperfect for every maximal ideal $\mathfrak{m}$ of $R$.
From now on we are using the conventions of Notation 5.1.
The following example explains a particular instance of the appearance of locally semiperfect algebras that, in addition, satisfy that they have a simple artinian ring of quotients. With the additional hypothesis that the domain $R$ is $h$-local, they will be our main object of study in Section 9.
@@
Proof. Let $I$ be a non-zero two-sided ideal of $\Lambda$ which is the trace of a projective right $\Lambda$-module. Since $\Lambda$ is torsion-free, $I_{\mathfrak{m}}\neq\{0\}$ holds for every maximal ideal $\mathfrak{m}$ of $R$. By Lemma 5.3 and Lemma 6.2, $I_{\mathfrak{m}}$ is also the trace of a finitely generated projective right $\Lambda_{\mathfrak{m}}$-module.
Let $S$ denote a simple right $\Lambda_{Q}$-module, and assume that $\Lambda_{Q}\cong S^{k}$ as a right $\Lambda_{Q}$-module. Since $\Lambda_{Q}$ is simple artinian and $I$ is non-zero, by Lemma 5.9, there exists a finite set $\mathcal{M}$ of maximal ideals of $R$ such that $I_{\mathfrak{m}}=\Lambda_{\mathfrak{m}}$ if $\mathfrak{m}\not\in\mathcal{M}$. For any maximal ideal $\mathfrak{m}\in\mathcal{M}$, let $P(\mathfrak{m})$ be a finitely generated projective right $\Lambda_{\mathfrak{m}}$-module with trace ideal $I_{\mathfrak{m}}$. Then $P(\mathfrak{m})_{Q}\cong S^{t_{\mathfrak{m}}}$ for a suitable positive integer $t_{\mathfrak{m}}\geq 1$. Let $t=\mathrm{lcm}\{t_{\mathfrak{m}}\mid\mathfrak{m}\in\mathcal{M}\}$, write $t=\ell_{\mathfrak{m}}\cdot t_{\mathfrak{m}}$, and let $P^{\prime}(\mathfrak{m}):=P(\mathfrak{m})^{k\ell_{\mathfrak{m}}}$ for each maximal ideal $\mathfrak{m}$ of $\mathcal{M}$. For each $\mathfrak{m}\notin\mathcal{M}$, take $P^{\prime}(\mathfrak{m})=\Lambda_{\mathfrak{m}}^{t}$. Note that $P^{\prime}(\mathfrak{m})_{Q}\cong S^{tk}\cong\Lambda_{Q}^{t}$ and $\operatorname{Tr}_{\Lambda_{\mathfrak{m}}}(P^{\prime}(\mathfrak{m}))=I_{\mathfrak{m}}$ hold for every maximal ideal $\mathfrak{m}$ of $R$. By Package Deal Theorem 5.13 applied to the package $\{P^{\prime}(\mathfrak{m})\mid\mathfrak{m}\in\operatorname{mSpec}(R)\}$, there exists a finitely generated projective right $\Lambda$-module $P$ such that $P_{\mathfrak{m}}\cong P^{\prime}(\mathfrak{m})$ for any maximal ideal $\mathfrak{m}$ of $R$. By Lemma 5.3, $\operatorname{Tr}_{\Lambda}(P)_{\mathfrak{m}}=\operatorname{Tr}_{\Lambda_{\mathfrak{m}}}(P^{\prime}(\mathfrak{m}))=I_{\mathfrak{m}}$ for any maximal ideal $\mathfrak{m}$ of $R$. This implies that $I=\operatorname{Tr}_{\Lambda}(P)+I=\operatorname{Tr}_{\Lambda}(P)$.
A similar argument gives the conclusion for the traces of projective left $\Lambda$-modules due to the symmetry of the hypotheses.
The equality $\mathcal{T}_{r}(\Lambda)=\mathcal{T}_{\ell}(\Lambda)$ follows from the canonical bijection between finitely generated projective right and left $\Lambda$-modules induced by the $\Lambda$-dual, and because the trace of a finitely generated projective module coincides with the trace of its $\Lambda$-dual, cf. Remark 2.6.
@@
Let $R$ be a commutative domain of finite character with field of fractions $Q$. Let $\Lambda$ be a locally semiperfect torsion-free $R$-algebra such that $\Lambda_{Q}$ is simple artinian; so that there exist $k$ and a division ring D such that $\Lambda_{Q}\cong M_{k}(D)$. Then, for almost all maximal ideals $\mathfrak{m}$ of $R$, the ring $\Lambda_{\mathfrak{m}}$ has the following properties:
@@
    There is exactly one indecomposable projective $\Lambda_{\mathfrak{m}}$-module up to isomorphism. Moreover, $\Lambda_{\mathfrak{m}}\cong M_{k}(T_{\mathfrak{m}})$, where $T_{\mathfrak{m}}$ is a local ring such that $T_{\mathfrak{m}}\otimes_{R}Q\cong D$.
@@
    If $U$ and $V$ are projective $\Lambda_{\mathfrak{m}}$-modules, then $U\cong V$ if and only if $U_{Q}\cong V_{Q}$.
Proof. Let $S$ be a simple $\Lambda_{Q}$-module and assume that $\Lambda_{Q}\cong S^{k}$. Let $X$ be a finitely generated $\Lambda$-module, which is torsion-free as an $R$-module, and such that $X_{Q}\cong S$. By Lemma 5.8, $X_{\mathfrak{m}}^{k}$ is free for almost all maximal ideals. If $\mathfrak{m}$ is such a maximal ideal, then $X_{\mathfrak{m}}^{k}\cong\Lambda_{\mathfrak{m}}$ gives an indecomposable decomposition of the semiperfect ring $\Lambda_{\mathfrak{m}}$. Hence $X_{\mathfrak{m}}$ is the only indecomposable projective $\Lambda_{\mathfrak{m}}$-module up to isomorphism. As $\Lambda_{\mathfrak{m}}$ is semiperfect, the endomorphism ring of $X_{\mathfrak{m}}$ is a local ring that we will denote by $T_{\mathfrak{m}}$. Hence $\Lambda_{\mathfrak{m}}\cong\mathrm{End}_{\Lambda_{\mathfrak{m}}}(X_{\mathfrak{m}}^{k})\cong M_{k}(T_{\mathfrak{m}}).$ By Lemma 5.2, $T_{\mathfrak{m}}\otimes_{R}Q\cong D$.
@@
Let $R$ be an $h$-local domain with field of fractions $Q$, and let $\Lambda$ be a locally semiperfect torsion-free $R$-algebra such that $\Lambda_{Q}$ is simple artinian, so that there exist $k$ and a division ring D such that $\Lambda_{Q}\cong M_{k}(D)$. In view of Lemma 6.9, the maximum spectrum of $R$ can be written as a disjoint union of two sets $A$ and $B$ such that for any $\mathfrak{m}\in A$, $\Lambda_{\mathfrak{m}}$ is a ring of $k\times k$ matrices over a local ring and $B$ is a finite, non-empty set. In particular, for $\mathfrak{m}\in A$, $\mathcal{T}_{r}(\Lambda_{\mathfrak{m}})=\mathcal{T}_{\ell}(\Lambda_{\mathfrak{m}})=\{0,\Lambda_{\mathfrak{m}}\}$.
In view of Package Deal Theorem 5.11 and Lemma 6.7, there is a bijective correspondence
@@
Let $R$ be an $h$-local domain with field of fractions $Q$, and let $\Lambda$ be a locally semiperfect torsion-free $R$-algebra such that $\Lambda_{Q}$ is simple artinian. Then $\mathcal{T}(\Lambda)$ is finite. Moreover, if $P$ is a finitely generated projective $\Lambda$-module, then $I=\operatorname{Tr}_{\Lambda}(P)$ is a minimal non-zero element of $\mathcal{T}(\Lambda)$ if and only if $P_{\mathfrak{m}}$ is a finite power of an indecomposable projective $\Lambda_{\mathfrak{m}}$-module for any $\mathfrak{m}\in\operatorname{mSpec}(R)$.
Proof. Finiteness of ${\mathcal{T}}(\Lambda)$ follows from Remark 6.10. Also, note that the correspondence of the remark preserves inclusions, so $\operatorname{Tr}_{\Lambda}(P)$ is a minimal element of ${\mathcal{T}}(\Lambda)\setminus\{0\}$ if and only if $\operatorname{Tr}_{\Lambda_{\mathfrak{m}}}(P_{\mathfrak{m}})$ is a minimal element of ${\mathcal{T}}(\Lambda_{\mathfrak{m}})\setminus\{0\}$ for any $\mathfrak{m}\in\operatorname{mSpec}(R)$ if and only if $P_{\mathfrak{m}}$ is a finite power of an indecomposable projective $\Lambda_{\mathfrak{m}}$-module.
@@
  1. (i)
    $P$ is finitely generated if and only if $P_{Q}$ is a finitely generated $\Lambda_{Q}$-module if and only if there exists a maximal ideal $\mathfrak{m}$ of $R$ such that $P_{\mathfrak{m}}$ is a finitely generated $\Lambda_{\mathfrak{m}}$-module.
  2. (ii)
    The set $\mathcal{I}(P)=\{J\unlhd\Lambda\mid P/PJ\mbox{ is finitely generated}\}$ has a minimal element, which is the trace of a finitely generated, projective right $\Lambda$-module. If $I$ is such ideal, then it is characterized by the property that, for any maximal ideal $\mathfrak{m}$ of $R$, $I_{\mathfrak{m}}$ is the minimal element of $\mathcal{I}(P_{\mathfrak{m}})=\{J\unlhd\Lambda\mid P_{\mathfrak{m}}/P_{\mathfrak{m}}J\mbox{ is finitely generated}\}$.
@@
$(i)$ We claim that if $P_{Q}\cong S^{n}$ for some $n\in\mathbb{N}$, then $P$ is finitely generated. Indeed, by hypothesis, there exists a finitely generated $\Lambda$-submodule $P_{0}$ of $P$ such that $(P_{0})_{Q}=P_{Q}$. Then we may assume that we have $P_{0}\leq P\leq\Lambda^{(\omega)}=P\oplus P^{\prime}$, and for any $p\in P$ there exists $0\neq d\in R$ such that $pd\in P_{0}$. Since $\Lambda$ is torsion-free as an $R$-module, this implies that there exists a finite subset $F$ of $\omega$ such that $P_{0}\leq P\subseteq\Lambda^{F}$. Hence, $\Lambda^{F}=P\oplus(P^{\prime}\cap\Lambda^{F})$ so $P$ is finitely generated.
This proves statement $(i)$.
$(ii)$ In the proof, we maintain the notation from Remark 6.10 regarding the partition of the maximal spectrum of $R$.
If $P$ is finitely generated, the statement is clear. Assume that $P$ is a countably generated, projective, right $\Lambda$-module that is not finitely generated. By $(1)$, $P_{Q}\cong S^{(\omega)}$. Therefore, there are two possible situations regarding the localizations of $P$ at the maximal ideals of $R$:
  1. (a)
    $P_{\mathfrak{m}}$ is a free $\Lambda_{\mathfrak{m}}$-module of infinite rank for any $\mathfrak{m}\in A$. Therefore, $\Lambda_{\mathfrak{m}}$ is the only element of the set $\mathcal{I}(P_{\mathfrak{m}})=\{J\unlhd\Lambda_{\mathfrak{m}}\mid P_{\mathfrak{m}}/P_{\mathfrak{m}}J\mbox{ is finitely generated}\}$.
  2. (b)
    $P_{\mathfrak{n}}$ is an infinitely generated projective module over the semiperfect ring $\Lambda_{\mathfrak{n}}$ for any $\mathfrak{n}\in B$. By Proposition 6.4, there exists a non-zero ideal $I(\mathfrak{n})$ that is the minimal element of the set $\mathcal{I}(P_{\mathfrak{n}})=\{J\unlhd\Lambda_{\mathfrak{n}}\mid P_{\mathfrak{n}}/P_{\mathfrak{n}}J\mbox{ is finitely generated}\}$. In addition, $I(\mathfrak{n})$ is the trace of a finitely generated projective right $\Lambda_{\mathfrak{n}}$-module.
By Package Deal Theorem 5.11, there exists an ideal $I$ of $\Lambda$ such that $I_{\mathfrak{m}}=\Lambda_{\mathfrak{m}}$ for any $\mathfrak{m}\in A$ and $I_{\mathfrak{n}}=I(\mathfrak{n})$ for any $\mathfrak{n}\in B$. In addition, $I$ is the trace of a (finitely generated, by Lemma 6.7) projective right $\Lambda$-module. We claim that $I$ is the minimal element of $\mathcal{I}(P)$.
@@
which is finitely generated. Therefore, $I\in\mathcal{I}(P)$.
Let $I^{\prime}\in\mathcal{I}(P)$ be such that $I^{\prime}\subseteq I$. Then, because $\Lambda$ is torsion-free as an $R$-module, for any maximal ideal $\mathfrak{m}$ of $R$, $I^{\prime}_{\mathfrak{m}}\in\mathcal{I}(P_{\mathfrak{m}})$ and $(I^{\prime})_{\mathfrak{m}}\subseteq I_{\mathfrak{m}}=I(\mathfrak{m})$. The minimality in the choice of $I(\mathfrak{m})$ in $(a)$ and $(b)$ implies that $(I^{\prime})_{\mathfrak{m}}=I_{\mathfrak{m}}=I(\mathfrak{m})$ for any maximal ideal $\mathfrak{m}$ of $R$. Hence, $I^{\prime}=I$. This shows that $I$ is the minimal element of $\mathcal{I}(P)$.
The proof also shows that for any maximal ideal $\mathfrak{m}$ of $R$, $I_{\mathfrak{m}}$ is the minimal element of $\mathcal{I}(P_{\mathfrak{m}})$.
@@
Let $R$ be a commutative domain with field of fractions $Q$, and let $\Lambda$ be a torsion-free $R$-algebra. Let $I\in\mathcal{T}(\Lambda)$. Let $P$ be a countably generated projective $\Lambda$-module, and let $V$ be a finitely generated $\Lambda$-module which is torsion-free as an $R$-module. Let $\mathfrak{p}$ be a prime ideal of $R$ such that $V_{\mathfrak{p}}=V_{\mathfrak{p}}I_{\mathfrak{p}}$ and $P_{\mathfrak{p}}=\bigoplus_{i\in\mathbb{N}}U_{i}$, where each $U_{i}$ is a finitely generated $\Lambda_{\mathfrak{p}}$-module, such that infinitely many of these $U_{i}$ are isomorphic to a $\Lambda_{\mathfrak{p}}$-module $U$ whose trace ideal contains $I_{\mathfrak{p}}$.
Then, for every finitely generated submodule $P_{1}$ of $P$, there exists a homomorphism $f\in\operatorname{Hom}_{\Lambda}(P,V)$ such that $f(P_{1})=0$, and $f_{\mathfrak{p}}$ is onto.